OpenStudy (anonymous):

Find the exact value for sin(x+y) if sinx = -4/5 and cosy = 15/17. Angles x and y are in the IV quadrant.

OpenStudy (watchmath):

Try to use the formula \(\sin(x+y)=\sin x\cos y +\cos x\sin y\)